# thanks: Corresponding author

### v3

@@ Abstract
Due to the success of the Standard Model (SM), it is reasonable to anticipate that the signal of new physics (NP) beyond the SM is small. Consequently, future searches for NP and precision tests of the SM will require high luminosity collider experiments. Moreover, as precision tests advance, rare processes with many final-state particles require consideration which demand the analysis of a vast number of observables. The high luminosity produces a large amount of experimental data spanning a large observable space, posing a significant data-processing challenge. In recent years, quantum machine learning has emerged as a promising approach for processing large amounts of complex data on a quantum computer. In this study, we propose quantum searching neighbor (QSN) and variational QSN (VQSN) algorithms to search for NP. The QSN is a classification algorithm. The VQSN introduces variation to the QSN to process classical data. As applications, we apply the (V)QSN in the phenomenological study of the NP at the Large Hadron Collider and muon colliders. The results indicate that the VQSN demonstrates superior efficiency to a classical counterpart k-nearest neighbor algorithm, even when dealing with classical data.
I Introduction
@@
The standard model (SM) is supported by a substantial body of experimental evidence and can be concluded to describe and explain the majority of phenomena in particle physics, with a few rare exceptions. These exceptions include experimental results such as the neutrino mass [1, 2, 3], the $W$ boson mass [4, 5], the muon anomalous magnetic moment ($g-2$[6, 7, 8], and other issues remain unresolved [9]. Furthermore, the SM is unable to account for the existence of dark matter and the force of gravity. Consequently, there is a widely held belief in the existence of new physics (NP) beyond the SM, and the search for NP and precision tests of the SM have become a prominent area of interest within the high-energy physics (HEP) community [10]. Given the success of the SM, it is reasonable to posit that any NP signals that have yet to be discovered will be exceedingly weak. Therefore, the search for NP and the precision tests of the SM motivate higher luminosity collider experiments. The processing of large amounts of data is inherent to such experiments, and thus places demands on our ability to process large amounts of data efficiently.
@@
The application of machine learning (ML) algorithms represents a promising avenue for the efficient processing of data. Its applications have already been demonstrated in HEP [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Of particular interest in the study of NP phenomenology is the employment of the anomaly detection (AD) algorithms, which have emerged as a prominent area of research in recent years [24, 25, 26, 27, 28, 29]. Since ML methods often eliminate the need for detailed kinematic analysis, their benefits become increasingly evident as kinematics grow more complex. Consequently, the advantages will become increasingly apparent as further advances in the precision tests of the SM will inevitably involve the study of rare processes with larger numbers of particles in the final states.
@@
In the meantime, quantum computing represents another effective approach for processing large volumes of data. Despite quantum computing currently being in the noisy intermediate-scale quantum (NISQ) [30, 31, 32] era, there has been a notable increase in research activity within the HEP community related to this field [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. Many ML algorithms can be implemented by quantum computing [53, 54, 55]. In conjunction with the advancement of quantum ML algorithms, there has been a growing interest in the deployment of quantum ML in the HEP phenomenological studies, encompassing techniques such as variational quantum classifier (VQC) [56, 57, 58, 59], quantum support vector machine (QSVM) [56, 60, 61, 62], and quantum kernel k-means (QKKM) [63]. In this study, we propose a variational quantum searching neighbor (VQSN) algorithm for identifying NP signals.
@@
To illustrate our algorithm, the SM effective field theory (SMEFT) [64, 65, 66, 67] are used as the test cases. The SMEFT has recently been employed extensively in the search for NP. When the collision energy is insufficient to excite NP particles, the NP particles can be integrated out, and the NP effects become new interactions of known particles. These interactions are formally represented as higher-dimensional operators with Wilson coefficients suppressed by powers of a NP scale, denoted as $\Lambda$. The signals of the operators that are most likely to be found correspond to those with Wilson coefficients least suppressed by $\Lambda$. Instead of dealing with various NP models, the number of operators to be considered at a specific order of $\Lambda$ is finite. Moreover, if an operator is not found, all NP models contributing to this operator can be constrained. Therefore, the use of the SMEFT is in line with the theme of high efficiency in this work. The present study will investigate the utility of VQSN for the search for high operators at the large hadron collider (LHC) and future muon colliders. VQSN demonstrates superior performance compared to a classical k-nearest neighbor (KNN) algorithm, even when applied to classical datasets.
II Searching for neighbors
In the search for NP, the data to be dealt with are the collision events. The idea of this work is to decide whether some unknown events are neighbors of the SM events, or the NP events. The neighbor events can be defined as soon as the distance between two events is defined. In the case of HEP, there are different methods to define the distance between two events, which can have an impact on the results of event selection strategies [68]. In this work, the observables of final-state particles in one collision event are mapped to a quantum state, and we use a quantum computing friendly way, i.e. the overlap between two states to define distance.
II.1 The circuit to search for neighbors
Figure 1: The circuits used in this paper. The top left panel shows the circuit to construct $\langle\phi|\psi_{n}\rangle|n\rangle$. The top right panel shows the circuit to fit the ansatz. The bottom left panel shows the circuit to find the classification assignment of $|\phi\rangle$. The bottom right panel shows an example of ‘hardware efficient’ ansatz used in this paper.
A circuit can be constructed to find the vectors with higher probabilities for closer neighbors. The first step is to construct $|\psi_{n}\rangle_{a}|n\rangle_{b}$ which can be realized using amplitude encoding, the subscript ‘a’ and ‘b’ denote the corresponding registers for clarity. When the $|\psi_{n}\rangle_{a}|n\rangle_{b}$ state is created, apply on register $a$ the inverse of the circuit to amplitude encode $|\phi\rangle_{a}$. And then deduct measurements on the $a$ register, when the outcome is $|0\rangle$, the state obtained is then $|0\rangle_{a}\sum_{n}\langle\phi|\psi_{n}\rangle|n\rangle$. The success rate (denoted as $r$) to create the state is $r=\sum\left|\langle\phi|\psi_{n}\rangle\right|^{2}/N$, where $N$ is the number of $\psi_{n}$ vectors. When $|0\rangle_{a}\sum_{n}\langle\phi|\psi_{n}\rangle|n\rangle_{b}$ is obtained, a measurement on $b$ register will result in $n$ with the probability proportional to $\left|\langle\phi|\psi_{n}\rangle\right|^{2}$, therefore, $|n\rangle$ will be obtained with higher probabilities if $\psi_{n}$ are closer to $\phi$. Denoting the amplitude encoding circuits as $U_{|\psi_{n}\rangle|n\rangle}$ and $U_{|\phi\rangle}$, i.e., $U_{|\psi_{n}\rangle|n\rangle}|0\rangle=|\psi_{n}\rangle|n\rangle$ and $U_{|\phi\rangle}|0\rangle=|\phi\rangle$, a summary of the circuit to create $\langle\phi|\psi_{n}\rangle|n\rangle$ is depicted in Fig. 1. (a). In this work, we use the uniform rotation gates [69] to implement the amplitude encoding.
@@ II.1 The circuit to search for neighbors
Without assumptions on the distribution of the vectors, there is no reason to prefer any result of the measurement on $a$ register. Therefore, if the vectors are $d$-dimensional complex vectors, $r\approx 1/d$, which is verified by using complex uniform random numbers. The success rate can be improved to $r\approx 1/\sqrt{d}$ by amplitude amplification, or the ‘QSearch’ algorithm  [70].
II.2 Fitting the training set
@@
In this work, $|\psi_{n}\rangle|n\rangle$ represent events that would be generated by a Monte Carlo (MC) simulation. Since at present, there is no quantum MC algorithm to generate events, one has to deal with a classical dataset. An improvement is to approximate the state $|\psi_{n}\rangle|n\rangle$ with a parameterized ansatz circuit.
@@
Since our goal is to classify, in the following, we assume that each $\psi$ vector carries a classification label and that the states are stored accordingly. Thus, the state to be amplitude encoded is $|\psi_{k,n}\rangle_{a}|k\rangle_{b_{1}}|n\rangle_{b_{2}}$, where $k$ indicates the class. Denoting the circuit of the ansatz as $U_{A}(\alpha_{i})$ where $\alpha_{i}$ are trainable parameters, the circuit of fitting an ansatz is shown in Fig. 1. (b), where the probability of the outcome $|0\rangle$ in the measurements is $\left|\langle\Psi(\alpha_{i})|\psi_{k,n}\rangle|k\rangle_{b_{1}}|n\rangle_{b_{2}}\right|^{2}$ with $U_{A}(\alpha_{i})|0\rangle=|\Psi(\alpha_{i})\rangle$. The training process adjusts the parameters $\alpha_{i}$ to maximize the probability of obtaining $|0\rangle$ in the measurements.
II.3 Variation quantum searching neighbor algorithm
@@
In the case of QSN the step-2 is omitted and the exact $\langle\phi|\psi_{k,n}\rangle|k\rangle_{b_{1}}|n\rangle_{b_{2}}$ is created. The circuit of VQSN is shown in Fig. 1. (c).
In essence, the VQSN is a variant of the VQC algorithms. However, it possesses two advantages. Firstly, conventional VQC can be seen as the quantum analogue of an artificial neural network, which is a ‘black box’ difficult to interpret. In contrast, the state to build in QSN has a discernible geometric interpretation, and the target of the training is guaranteed to exist. Secondly, VQSN is accompanied by a classical algorithm, KNN, which allows for a comparison between quantum and classical algorithms. This is clearer for a binary classification problem, in which QSN can be viewed as using $\sum w_{n}t_{n}/N$ to determine the classification of $\phi$ where $w_{n}=\left|\langle\phi|\psi_{n}\rangle\right|^{2}$, and $t_{n}=0,1$ is the class of $\psi_{n}$. Therefore, it uses a weighted average of the entire training dataset. KNN can also be viewed as a weighted average except that $\{w_{n}=1|\psi_{n}\in S\}$ and $\{w_{n}=0|\psi_{n}\notin S\}$, where $S$ is the set contains the $k$ nearest $\psi_{n}$ vectors to $\phi$.
II.4 A numerical example
Before moving on to the search for NP, a numerical example is implemented to illustrate the algorithm. We use the ‘hardware efficient’ ansatz [71, 72] with the pairwise entanglement layer,
$\begin{split}U_{A}(\alpha_{i})&=\prod_{i}^{l}\left(\prod_{q}e^{{\rm i}\rho^{(q)}_{i}\sigma^{(q)}_{z}}\prod_{q}e^{{\rm i}\beta^{(q)}_{i}\sigma^{(q)}_{y}}\prod_{\langle q_{1},q_{2}\rangle}c_{z}^{q_{1},q_{2}}\right)\\ &\times\prod_{q}e^{{\rm i}\rho^{(q)}_{0}\sigma^{(q)}_{z}}\prod_{q}e^{{\rm i}\beta^{(q)}_{0}\sigma^{(q)}_{y}},\end{split}$ (1)
where $\sigma^{q}_{y,z}$ are Pauli matrices acting on $q$ qubit, the product over $q$ is the product over all qubits, $c_{z}^{q_{1},q_{2}}$ is the controlled-Z gate acting on $q_{1,2}$ qubits with $c_{z}=diag\left(1,1,1,-1\right)$, and $\langle q_{1},q_{2}\rangle$ is the set of pairs of qubits depending on the type of entanglement layer, $l$ is the number of layers of the ansatz, $\{\alpha_{i}\}=\{\beta^{(q)}_{0},\rho^{(q)}_{0},\beta^{(q)}_{1},\rho^{(q)}_{1},\ldots\beta^{(q)}_{l},\rho^{(q)}_{l}\}$ is a set of trainable parameters. The number of parameters is $2(l+1)n_{q}$ where $n_{q}$ is the number of qubits. $U_{A}$ is shown in Fig. 1. (d).

### v4

# A quantum machine learning classifier to search for new physics

@@ Abstract
Due to the success of the Standard Model (SM), it is reasonable to anticipate that the signal of new physics (NP) beyond the SM is small. Consequently, future searches for NP and precision tests of the SM will require high luminosity collider experiments. Moreover, as precision tests advance, rare processes with many final-state particles require consideration which demand the analysis of a vast number of observables. The high luminosity produces a large amount of experimental data spanning a large observable space, posing a significant data-processing challenge. In recent years, quantum machine learning has emerged as a promising approach for processing large amounts of complex data on a quantum computer. In this study, we propose quantum searching neighbor (QSN) and variational QSN (VQSN) algorithms to search for NP. The QSN is a classification algorithm. The VQSN introduces variation to the QSN to process classical data. As applications, we apply the (V)QSN in the phenomenological study of the NP at the Large Hadron Collider and muon colliders. Examples are implemented on a real quantum hardware, which confirms reliable performance under noisy conditions. The results indicate that the VQSN demonstrates superior efficiency in the sense of computational complexity to a classical counterpart k-nearest neighbor algorithm, even when dealing with classical data.
1 Introduction
@@
The standard model (SM) is supported by a substantial body of experimental evidence and can be concluded to describe and explain the majority of phenomena in particle physics, with a few rare exceptions. These exceptions include experimental results such as the neutrino mass Farzan:2017xzy ; Proceedings:2019qno ; Arguelles:2022tki , the WW boson mass CDF:2022hxs ; deBlas:2022hdk , the muon anomalous magnetic moment (g2g-2Muong-2:2021ojo ; Muong-2:2023cdq ; Aoyama:2020ynm , and other issues remain unresolved Crivellin:2023zui . Furthermore, the SM is unable to account for the existence of dark matter and the force of gravity. Consequently, there is a widely held belief in the existence of new physics (NP) beyond the SM, and the search for NP and precision tests of the SM have become a prominent area of interest within the high-energy physics (HEP) community Ellis:2012zz . Given the success of the SM, it is reasonable to posit that any NP signals that have yet to be discovered will be exceedingly weak. Therefore, the search for NP and the precision tests of the SM motivate higher luminosity collider experiments. The processing of large amounts of data is inherent to such experiments, and thus places demands on our ability to process large amounts of data efficiently.
@@
The application of machine learning (ML) algorithms represents a promising avenue for the efficient processing of data. Its applications have already been demonstrated in HEP Innocente:1992gq ; Albertsson:2018maf ; Guest:2018yhq ; Radovic:2018dip ; Baldi:2014kfa ; Ren:2017ymm ; Abdughani:2018wrw ; Ren:2019xhp ; Letizia:2022xbe ; DAgnolo:2019vbw ; DAgnolo:2018cun ; DeSimone:2018efk ; Lee:2018xtt . Of particular interest in the study of NP phenomenology is the employment of the anomaly detection (AD) algorithms, which have emerged as a prominent area of research in recent years MdAli:2020yzb ; Fol:2020tva ; Kasieczka:2021xcg ; CrispimRomao:2020ucc ; vanBeekveld:2020txa ; Kuusela:2011aa . Since ML methods often eliminate the need for detailed kinematic analysis, their benefits become increasingly evident as kinematics grow more complex. Consequently, the advantages will become increasingly apparent as further advances in the precision tests of the SM will inevitably involve the study of rare processes with larger numbers of particles in the final states.
@@
In the meantime, quantum computing represents another effective approach for processing large volumes of data. Despite quantum computing currently being in the noisy intermediate-scale quantum (NISQ) Chen:2021num ; Arute:2019zxq ; Preskill:2018jim era, there has been a notable increase in research activity within the HEP community related to this field Feynman:1981tf ; Roggero:2018hrn ; Roggero:2019myu ; Bauer:2022hpo ; Carena:2022kpg ; Gustafson:2022xdt ; Lamm:2024jnl ; Carena:2024dzu ; Atas:2021ext ; Li:2023vwx ; Cui:2019sfz ; Zou:2021pvl ; Georgescu:2013oza ; Lamm:2019uyc ; Li:2021kcs ; Echevarria:2020wct ; Perez-Salinas:2020nem ; Jordan:2011ci ; Mueller:2019qqj ; Zhu:2024own . Many ML algorithms can be implemented by quantum computing Biamonte:2016ugo ; qml2 ; Garcia:2022cqq . In conjunction with the advancement of quantum ML algorithms, there has been a growing interest in the deployment of quantum ML in the HEP phenomenological studies, encompassing techniques such as variational quantum classifier (VQC) Guan:2020bdl ; Wu:2020cye ; Grant:2018oml ; Terashi:2020wfi , quantum support vector machine (QSVM) Guan:2020bdl ; Wu:2021xsj ; Zhang:2023ykh ; Fadol:2022umw , and quantum kernel k-means (QKKM) Zhang:2024ebl . In this study, we propose a variational quantum searching neighbor (VQSN) algorithm for identifying NP signals.
@@
To illustrate our algorithm, the SM effective field theory (SMEFT) Weinberg:1979sa ; Grzadkowski:2010es ; Willenbrock:2014bja ; Masso:2014xra are used as the test cases. The SMEFT has recently been employed extensively in the search for NP. When the collision energy is insufficient to excite NP particles, the NP particles can be integrated out, and the NP effects become new interactions of known particles. These interactions are formally represented as higher-dimensional operators with Wilson coefficients suppressed by powers of a NP scale, denoted as $\Lambda$. The signals of the operators that are most likely to be found correspond to those with Wilson coefficients least suppressed by $\Lambda$. Instead of dealing with various NP models, the number of operators to be considered at a specific order of $\Lambda$ is finite. Moreover, if an operator is not found, all NP models contributing to this operator can be constrained. Therefore, the use of the SMEFT is in line with the theme of high efficiency in this work. The present study will investigate the utility of VQSN for the search for high operators at the large hadron collider (LHC) and future muon colliders. VQSN demonstrates superior performance compared to a classical k-nearest neighbor (KNN) algorithm, even when applied to classical datasets.
2 Searching for neighbors
2.1 The quantum feature space
Figure 1: The problem addressed by (V)QSN. Given known classifications of some points in a feature space (e.g., $32$ colored points in a $2$-dimensional space as shown in the figure), how to classify an unknown point (marked as $\times$’ in the figure). (V)QSN classifies $\times$’ based on its distance to the known points, assigning it to the class belonging to the known points closer to it (its neighbors).
(V)QSN is a classification method. In the next section, we will explain why a classification method is used to search for NP, but in this section, we will focus on the classification method itself. The problem (V)QSN addresses is, if we know some points and their classifications, how should unknown points be classified? Fig. 1 provides an example. In Fig. 1, $32$ known points are classified into four categories (indicated by different colors). These $32$ points with known classification assignment constitute the training set and are generated using scikit-learn Pedregosa:2011ork , and rescaled to range $1<x_{i}<1$ and $1<y_{i}<1$ where $x_{i}$ and $y_{i}$ are the coordinates of the points. How should the unknown point (marked with an $\times$’) be classified? V(QSN) classifies points based on the distance between known points and unknown points. If an unknown point is closer to the known points of a certain class, it will be classified into that category.
To use quantum computing for finding neighbors, the first step is to map the data points to quantum states. There are many schemes capable of establishing this one-to-one mapping. While these different schemes will impact the algorithm’s performance, they do not affect its implementation. Taking Fig. 1 as an example, a straightforward way is to map these two-dimensional points to the points on the Bloch sphere. For simplicity, $x$, $y$ coordinates are translated as zenith and azimuth angles of the Bloch sphere, i.e., for the point $\left(x_{i},y_{i}\right)$ the corresponding quantum state is $|\phi_{i}\rangle=e^{i(y_{i}+1)\pi/2}\cos\left((x_{i}+1)\pi/4\right)|0\rangle+e^{-i(y_{i}+1)\pi/2}\sin\left((x_{i}+1)\pi/4\right)|1\rangle$.
2.2 The circuit to search for neighbors
Figure 2: The circuit to construct the state $\langle\phi|\psi_{n}\rangle|n\rangle$. $U_{|\psi_{j}\rangle_{a}|j\rangle_{b}}$ and $U_{|\phi\rangle}$ are the circuits to encode the training set points $U_{|\psi_{j}\rangle_{a}|j\rangle_{b}}|0\rangle_{a}|0\rangle_{b}=|\psi_{j}\rangle_{a}|j\rangle_{b}$ and test point $U_{|\phi\rangle}|0\rangle=|\phi\rangle$, respectively. If the measurement on $a$ register yields the $|0\ldots 0\rangle_{a}$ state, the result is $|0\ldots 0\rangle_{a}\langle\phi|\psi_{j}\rangle|j\rangle_{b}$.
Since the concept of distance is used for classification, a crucial step is defining distance which is convenient to calculate in the context of quantum computing. After establishing a one-to-one correspondence between data points and quantum states, a distance measure suitable for quantum computing is the overlap between two states. For example, one can define $d_{ij}=\sqrt{1-\left|\langle\phi_{i}|\psi_{j}\rangle\right|^{2}}$, where $i,j$ denote two data points, $\phi_{i}$ and $\psi_{j}$ represent their corresponding quantum states. In this way, the task of finding neighbors for data point $i$ translates into finding those $\psi$ states which have a larger overlap with the state $\phi_{i}$.
The overlap between two states can be calculated using the control-swap test. However, as will be shown, one can leverage the superposition property of quantum states to directly compute the overlaps of multiple $\psi_{j}$ states with a state $\phi$, storing the results as amplitude coefficients. In this way, when measuring the quantum state where the probability of obtaining a specific outcome is proportional to the square of the absolute value of its amplitude coefficient, states with larger overlaps are naturally measured with higher probability. This allows us to identify the neighbors of $\phi$.
As mentioned above, the problem becomes: given one $\phi$ and multiple $\psi_{j}$, how does one obtain the state $\langle\phi|\psi_{j}\rangle|j\rangle$. Assume $U_{|\phi\rangle}$ is the circuit that creates the quantum state $\phi$, i.e., $U_{|\phi\rangle}|0\rangle=|\phi\rangle$. For ease of description, we divide the registers into $a$ and $b$, where $a$ corresponds to the qubits storing $\psi_{j}$ and $b$ corresponds to the indices. For the state $|\psi_{j}\rangle_{a}|j\rangle_{b}$, if we apply the inverse of $U_{|\phi\rangle}$ to register $a$ and measuring the $a$ register yields the $|0\ldots 0\rangle_{a}$ state, then the state becomes $|0\ldots 0\rangle_{a}\langle\phi|\psi_{j}\rangle|j\rangle_{b}$. Note that the state corresponding to register $b$ at this point is precisely what we need. Denoting $U_{|\psi_{j}\rangle_{a}|j\rangle_{b}}$ as the circuit to create the state $|\psi_{j}\rangle_{a}|j\rangle_{b}$, i.e., $U_{|\psi_{j}\rangle_{a}|j\rangle_{b}}|0\rangle_{a}|0\rangle_{b}=|\psi_{j}\rangle_{a}|j\rangle_{b}$, the circuit is shown in Fig. 2. In this work, we use the uniform rotation gates Mottonen:2004vly to create the state $|\psi_{j}\rangle_{a}|j\rangle_{b}$
@@ 2.2 The circuit to search for neighbors
The success rate (denoted as $r$) to create the state $\langle\phi|\psi_{j}\rangle|j\rangle$ (the probabilities that $|0\ldots 0\rangle_{a}$ is measured) is $r=\sum\left|\langle\phi|\psi_{n}\rangle\right|^{2}/N$, where $N$ is the number of $\psi_{n}$ vectors. Without assumptions on the distribution of the vectors, there is no reason to prefer any result of the measurement on $a$ register. Therefore, if the vectors are $d$-dimensional complex vectors, $r\approx 1/d$, which is verified by using complex uniform random numbers. The success rate can be improved to $r\approx 1/\sqrt{d}$ by amplitude amplification, or the ‘QSearch’ algorithm  Brassard:2000xvp .
2.3 Fitting the training set
Figure 3: The circuit to fit the ansatz. $U_{A}$ denotes the ansatz circuit with trainable parameters $\alpha_{i}$. The goal is to adjust $\alpha_{i}$ to maximize the probability where $|0\ldots 0\rangle$ is measured. An example of $U_{A}$ with the ‘SCA’ layer is shown.
@@ 2.3 Fitting the training set
The complexity of finding neighbor algorithm is mainly the query complexity. In this work, $|\psi_{n}\rangle|n\rangle$ represent events that would be generated by a Monte Carlo (MC) simulation. Since at present, there is no quantum MC algorithm to generate events, one has to deal with a classical dataset. An improvement is to approximate the state $|\psi_{n}\rangle|n\rangle$ with a parameterized ansatz circuit.
@@
Denoting the circuit of the ansatz as $U_{A}(\alpha_{i})$ where $\alpha_{i}$ are trainable parameters, the circuit of fitting an ansatz is shown in Fig. 3, where the probability of the outcome $|0\ldots 0\rangle$ in the measurements is $\left|\langle\Psi(\alpha_{i})|\psi_{j}\rangle_{a}|j\rangle_{b}\right|^{2}$ with $U_{A}(\alpha_{i})|0\rangle=|\Psi(\alpha_{i})\rangle$. The training process adjusts the parameters $\alpha_{i}$ to maximize the probability of obtaining $|0\ldots 0\rangle$ in the measurements.
In this work, we use the ‘hardware efficient’ ansatz Kandala:2017vok ; Bravo-Prieto:2019kld with entanglement layers,
[…]
where $\sigma^{q}_{y,z}$ are Pauli matrices acting on $q$ qubit, the product over $q$ is the product over all qubits, $c_{z}^{q_{1},q_{2}}$ is the controlled-Z gate acting on $q_{1,2}$ qubits with $c_{z}=diag\left(1,1,1,-1\right)$, and $\langle q_{1},q_{2}\rangle$ is the set of pairs of qubits depending on the type of entanglement layer, $l$ is the number of layers of the ansatz, $\{\alpha_{i}\}=\{\beta^{(q)}_{0},\rho^{(q)}_{0},\beta^{(q)}_{1},\rho^{(q)}_{1},\ldots\beta^{(q)}_{l},\rho^{(q)}_{l}\}$ is a set of trainable parameters. The number of parameters is $2(l+1)n_{q}$ where $n_{q}$ is the number of qubits. Two different entanglement layers are used in this work, which are the ‘shifted-circular-alternating’ (SCA) layer Sim:2019yyv and the ‘pairwise’ layer. An example of $U_{A}$ with the ‘SCA’ layer is shown in Fig. 3. The ‘pairwise’ layer is shown in Sec. 2.5.
The relationship between training cost and training set is not straightforward to determine, as it depends on multiple factors: the expressive power of the ansatz, the gradient descent algorithm used during training, the distribution of the training dataset, and the method of encoding classical data. In the paper, the ansatz, gradient descent algorithm, and classical data encoding scheme were chosen naively, each of which admits further optimization, a direction we defer to future work. A key point is that, as with most machine learning algorithms, the training dataset plays a critical role. However, the distribution of the training data depends on the case to be studied, making it difficult to predict a priori whether or to what extent additional training cost would ensure convergence. One can, however, roughly estimate its scalability. Since the information content of the classical data is proportional to $Nd$, the cost of preparing the training dataset is at least $\mathcal{O}(Nd)$, regardless of the encoding method. Accordingly, the training cost is roughly on the order of the number of convergence iterations multiplied by $Nd$.
2.4 Variation quantum searching neighbor algorithm
Figure 4: The circuit to find the classification assignment of $|\phi\rangle$. If the measurement on the $a$ register yields $|0\ldots 0\rangle$, then $b_{1}$ is measured. If the quantum states corresponding to the points classified as $k$ in the training set have larger overlaps with $|\phi\rangle$, the measurement on $b_{1}$ yields outcome $|k\rangle_{b_{1}}$ with higher probability.
Since our goal is to classify, in the following, we assume that each $\psi$ vector carries a classification label and that the states are stored accordingly. Thus, the state to be create is $|\psi_{k,n}\rangle_{a}|k\rangle_{b_{1}}|n\rangle_{b_{2}}$, where $k$ indicates the class, the $b$ register is split to $b_{1}$ and $b_{2}$ for clarity. With this ordering, it becomes unnecessary to measure the entire $b$ register to find neighbors of $\phi$. Instead, only measuring $b_{1}$ reveals which class $\phi$’s neighbors belong to.
@@ 2.4 Variation quantum searching neighbor algorithm
In the case of QSN the step-2 is omitted and the exact $\langle\phi|\psi_{k,n}\rangle|k\rangle_{b_{1}}|n\rangle_{b_{2}}$ is created. The circuit of VQSN is shown in Fig. 4.
In essence, the VQSN is a variant of the VQC algorithms. As a variant of VQC, given ansatze of comparable complexity their computational complexities are of the same order of magnitude. However, VQSN possesses two advantages. Firstly, conventional VQC can be seen as the quantum analogue of an artificial neural network, which is a ‘black box’ difficult to interpret. The reason is that VQC trains a parameterized quantum circuit to achieve the mapping $f(\psi_{sig})=1$ and $f(\psi_{bg})=0$, where $\phi_{sig}$ and $\phi_{bg}$ refer to the signal and background, respectively. This parameterized quantum circuit is constructed as a layered, interconnected ansatz, structurally resembling a neural network. The VQSN also trains a parameterized quantum circuit to realize the same discriminative effect. In contrast, the state to build in VQSN has a discernible geometric interpretation, and the target of the training is guaranteed to exist. Secondly, QSN is accompanied by a classical algorithm, KNN, which allows for a comparison between quantum and classical algorithms. This is clearer for a binary classification problem, in which QSN can be viewed as using $\sum w_{n}t_{n}/N$ to determine the classification of $\phi$ where $w_{n}=\left|\langle\phi|\psi_{n}\rangle\right|^{2}$, and $t_{n}=0,1$ is the class of $\psi_{n}$. Therefore, it uses a weighted average of the entire training dataset. KNN can also be viewed as a weighted average except that $\{w_{n}=1|\psi_{n}\in S\}$ and $\{w_{n}=0|\psi_{n}\notin S\}$, where $S$ is the set contains the $k$ nearest $\psi_{n}$ vectors to $\phi$.
Although our approach resembles KNN, it is fundamentally distinct. Unlike Refs. Lloyd:2013lby ; Wiebe:2014imz , we define distance through quantum overlap (same as Ref. basheer2021quantumknearestneighborsalgorithm ). This enables efficient identification of classification assignments via measurement. While QSN could extend to quantum KNN using the maxima finding algorithm from Ref. basheer2021quantumknearestneighborsalgorithm , VQSN deliberately eliminates this step to reduce quantum gate count and circuit depth. Furthermore, replacing data queries with an ansatz circuit significantly decreases gate requirements and depth, enabling execution of VQSN on a current NISQ hardware.
2.5 The numerical result
Before moving on to the search for NP, a numerical example is implemented to illustrate the algorithm.